# Transitional Uncertainty with Layered Intermediate Predictions

### v1

@@ Abstract
In this paper, we discuss feature engineering for single-pass uncertainty estimation. For accurate uncertainty estimates, neural networks must extract differences in the feature space that quantify uncertainty. This could be achieved by current single-pass approaches that maintain feature distances between data points as they traverse the network. While initial results are promising, maintaining feature distances within the network representations frequently inhibits information compression and opposes the learning objective. We study this effect theoretically and empirically to arrive at a simple conclusion: preserving feature distances in the output is beneficial when the preserved features contribute to learning the label distribution and act in opposition otherwise. We then propose Transitional Uncertainty with Layered Intermediate Predictions (TULIP) as a simple approach to address the shortcomings of current single-pass estimators. Specifically, we implement feature preservation by extracting features from intermediate representations before information is collapsed by subsequent layers. We refer to the underlying preservation mechanism as transitional feature preservation. We show that TULIP matches or outperforms current single-pass methods on standard benchmarks and in practical settings where these methods are less reliable (imbalances, complex architectures, medical modalities). Code is available at https://github.com/olivesgatech/TULIP
@@ 1 Introduction
Despite their critical importance, preserving features is not trivial in neural networks. In particular, information compression is a desirable property of neural networks and a central component of the learning problem (Tishby et al., 2000). In spite of this discrepancy, current single-pass uncertainty methods preserve features by maintaining distances between data points in the output and risk inhibiting compression of application irrelevant information (Liu et al., 2020; Van Amersfoort et al., 2020; van Amersfoort et al., 2021; Mukhoti et al., 2023; Kwon et al., 2020; Prabhushankar & AlRegib, ). As a result, several recent studies have shown several practical limitations of current single-pass methods such as their susceptibility to distributional shift (Postels et al., 2022). We provide a simple illustration of this effect in Figure 1a. The left plot shows the 2D neural network output features of two clusters when trained without explicit feature preservation. The network collapses the class clusters to single points creating a challenging setting for uncertainty estimation. The center plot shows the same 2D classification problem, but depicts a network trained with feature preservation constraints on the output. These disadvantages motivate the search for alternative feature preservation approaches.
@@
To this end, a more conservative strategy preserves features in the collection of multiple representations of the same sample (Malinin & Gales, 2018). From a preservation perspective, the approach is preferred as feature distances are encoded in differences (e.g. spread) of the individual sample representations and does not require constraining the compression property. We term this approach as transitional feature preservation and illustrate a toy example in the right plot of Figure 1a. We show three different representations of the same two clusters collected from different sources $h_{w_{1,2,3}}$. While the sources collapse points individually, the inter-source difference for a given sample reflects uncertainty (signified by $\Delta h$). A classic candidate that could implement this property are deep ensembles (Lakshminarayanan et al., 2017), where each source representation is collected from a separate network. However, ensembles face two central challenges for uncertainty estimation: 1) the evaluation requires several forward passes limiting their practicality; and 2) there is no guarantee that ensemble transitions preserve features as all ensemble members can collapse features to the same point collectively. We show an example of this phenomenon with a synthetic 2D spiral dataset (Figure 1b left plot). Uncertainties collected from ensemble transitions fail to capture uncertainty far away from the training data even though multiple networks are used to preserve features. Our method addresses both of these limitations. We provide experimental details and further discussion of our synthetic example in Appendix B.6 and C.2 respectively.
@@
In this paper, we address shortcomings of current single-pass uncertainty estimators and ensembles by leveraging intermediate representations. Figure 1b (right plot) shows the intuition behind our approach. Features are extracted from intermediate layers before information is collapsed allowing the uncertainty calculation within a single forward pass. From the uncertainty plot, we can further see the advantage of our approach in comparison to ensembles as uncertainty increases systematically further away from the training data showing the favorable properties of our approach. In addition to empirical evidence, we further provide theoretical guarantees for intermediate representations in Section 4. In particular, we show that the linear combination of representations is feature preserving when the first layer is collapse resistant. Finally, we combine intermediate representations with a single-pass uncertainty estimation layer, approximate Gaussian Processes, and arrive at a new single-pass model: Transitional Uncertainty with Layered Intermediate Predictions or TULIP in short. Our estimator requires less labeled training data and outperforms current single-pass estimators both on standard benchmarks and other data modalities (CT scans) with little additional space overhead. Further, we show that TULIP is preferable under complex architectures, class imbalance, and several other challenging settings for current single-pass estimators.
@@ 3.2 Distance Preservation under Class Imbalance
We compare the classification accuracy with and without spectral normalization in Figure 2. In addition to class imbalance, we further consider settings where we randomly remove training samples (left graph). We show this setting to determine that the accuracy gain/loss from output feature preservation is dependent on the available information, not the number of samples. Our experiments highlight both advantages and disadvantages of preserving features in the output representation: if the target distribution is sufficiently similar to the input distribution, additional preserved features correlate with the generalization objective and results in a performance increase. This can be seen from the accuracy improvement with spectral normalization under low imbalance severities or when samples are removed randomly. The opposite can be observed where, in contrast to random sample removal (left graph), we explicitly remove training and test samples to imbalance classes (right graph). Here, the training set contains more significant amounts of irrelevant information and spectral normalization significantly decreases the generalization performance.
@@ 4.2 Algorithm
Our method consists of three components: a principal feed-forward network, shallow-deep network exits with individual internal classifiers, and a combination head (Figure 3). During training, the shallow-deep network exits are trained jointly with the feed-forward component, while the combination head is fitted after optimization on a validation set extracted from the training data $\mathcal{X}_{ID}$. We emphasize that our method does not require out-of-distribution validation samples as the combination head is fitted with data from the training set. During inference, the input sample traverses both the feed-forward network, as well as the shallow exits. The final output prediction is derived from the main network, while the uncertainty score is derived from a combination of the intermediate output logits. In addition to our description in the main paper, we provide additional implementation details and algorithm pseudo code in Appendix D.
@@
Shallow-Deep networks (SDNs) (Kaya et al., 2019) were originally introduced in the context of computation reduction. A SDN is a modified version of a conventional DNN where additional internal classifiers are placed on intermediate representations to produce preliminary predictions. In our context, we utilize internal classifiers to produce intermediate logits for uncertainty estimation. Formally, given the intermediate layer $l$ of the principal feed-forward network with the latent representation $h^{l}_{w_{l}}(\mathbf{x})\in\mathbb{R}^{M_{l}}$, an internal classifier is a shallow network $f^{l}_{w_{l}}:\mathbb{R}^{M_{l}}\rightarrow\mathbb{R}^{K}$ that produces prediction logits for $K$ classes. Regarding positioning, we place internal classifiers uniformly across the feed forward network, similar to (Kaya et al., 2019).
@@ 7 Discussion and Limitations
A central observation we made in this work is that enforcing feature preservation by constraining model representations can be harmful to the model performance and highlight application relevance as a key requirement for effective representation constraints. In practice, the characteristic can be undesirable as training distributions can severely differ from deployment. We propose single-pass transitional feature preservation through intermediate representations to address these disadvantages. While our approach is effective, we do not claim that our improvements solve feature preservation in uncertainty estimation entirely. In particular, we propose one instance of TFP through SDNs that comes with its own set of limitations: similar to iterative methods, the success depends on the amount of source representations in $\Delta h$ to preserve features. For this purpose, SDNs are less effective on shallow architectures with fewer intermediate options to extract from. Further, we acknowledge that implementing internal classifiers comes with a large set of hyperparameters increasing space complexity. In our implementation, we use standard settings that show promising performance without significantly impacting the space complexity. However, other applications may require more elaborate hyperparameter explorations and additional internal classifiers. For this reason, we see reducing the space complexity of internal classifiers or the development of efficient hyperparameter tuners as promising research directions. Finally, we chose the combination of SDNs with Gaussian Processes out of simplicity and the strong empirical performance. However, key novelties of this paper (SDNs and singe-pass TFP) are not limited to one uncertainty score (Gaussian Processes) and are building blocks for single-pass uncertainty methods. We encourage researchers to implement different combinations of single-pass TFP and uncertainty scores to advance the field of single-pass uncertainty estimation.

### v2

@@ Abstract
In this paper, we discuss feature engineering for single-pass uncertainty estimation. For accurate uncertainty estimates, neural networks must extract differences in the feature space that quantify uncertainty. This could be achieved by current single-pass approaches that maintain feature distances between data points as they traverse the network. While initial results are promising, maintaining feature distances within the network representations frequently inhibits information compression and opposes the learning objective. We study this effect theoretically and empirically to arrive at a simple conclusion: preserving feature distances in the output is beneficial when the preserved features contribute to learning the label distribution and act in opposition otherwise. We then propose Transitional Uncertainty with Layered Intermediate Predictions (TULIP) as a simple approach to address the shortcomings of current single-pass estimators. Specifically, we implement feature preservation by extracting features from intermediate representations before information is collapsed by subsequent layers. We refer to the underlying preservation mechanism as transitional feature preservation. We show that TULIP matches or outperforms current single-pass methods on standard benchmarks and in practical settings where these methods are less reliable (imbalances, complex architectures, medical modalities).
@@ 1 Introduction
Despite their critical importance, preserving features is not trivial in neural networks. In particular, information compression is a desirable property of neural networks and a central component of the learning problem (Tishby et al., 2000). In spite of this discrepancy, current single-pass uncertainty methods preserve features by maintaining distances between data points in the output and risk inhibiting compression of application irrelevant information (Liu et al., 2020; Van Amersfoort et al., 2020; van Amersfoort et al., 2021; Mukhoti et al., 2023; Kwon et al., 2020; Prabhushankar & AlRegib, ). As a result, several recent studies have shown practical limitations of current single-pass methods such as their susceptibility to distributional shift (Postels et al., 2022). We provide a simple illustration of this effect in Figure 1a. The left plot shows the 2D neural network output features of two clusters when trained without explicit feature preservation. The network collapses the class clusters to single points creating a challenging setting for uncertainty estimation. The center plot shows the same 2D classification problem, but depicts a network trained with feature preservation constraints on the output. These disadvantages motivate the search for alternative feature preservation approaches.
@@
To this end, a more conservative strategy preserves features in the collection of multiple representations of the same sample (Malinin & Gales, 2018). From a preservation perspective, the approach is preferred as feature distances are encoded in differences (e.g. spread) of the individual sample representations and does not require constraining the compression property. We term this approach as transitional feature preservation and illustrate a toy example in the right plot of Figure 1a. We show three different representations of the same two clusters collected from different sources $h_{w_{1,2,3}}$. While the sources collapse points individually, the inter-source difference for a given sample reflects uncertainty (signified by $\Delta h$). A classic candidate that could implement this property are deep ensembles (Lakshminarayanan et al., 2017), where each source representation is collected from a separate network. However, ensembles face two central challenges for uncertainty estimation: 1) the evaluation requires several forward passes limiting their practicality; and 2) there is no guarantee that ensemble transitions preserve features because all ensemble members can collapse features to the same point collectively. We show an example of this phenomenon with a synthetic 2D spiral dataset (Figure 1b left plot). Uncertainties collected from ensemble transitions fail to capture uncertainty far away from the training data even though multiple networks are used to preserve features. Our method addresses both of these limitations. We provide experimental details and further discussion of our synthetic example in Appendix B.6 and C.2 respectively.
@@
In this paper, we address shortcomings of current single-pass uncertainty estimators and ensembles by leveraging intermediate representations. Figure 1b (right plot) shows the intuition behind our approach. Features are extracted from intermediate layers before information is collapsed allowing the uncertainty calculation within a single forward pass. From the uncertainty plot, we can see the advantage of our approach in comparison to ensembles. Uncertainty increases systematically further away from the training data showing while ensembles are uncertain only around decision boundaries. In addition to empirical evidence, we provide theoretical guarantees for intermediate representations in Section 4. In particular, we show that the linear combination of representations is feature preserving when the first layer is collapse resistant. Finally, we combine intermediate representations with a single-pass uncertainty estimation layer, approximate Gaussian Processes, and arrive at a new single-pass model: Transitional Uncertainty with Layered Intermediate Predictions or TULIP in short. Our estimator requires less labeled training data and outperforms current single-pass estimators both on standard benchmarks and other data modalities (CT scans) with little additional space overhead. Further, we show that TULIP is preferable under complex architectures, class imbalance, and several other challenging settings for current single-pass estimators.
@@ 3.2 Distance Preservation under Class Imbalance
We compare the classification accuracy with and without spectral normalization in Figure 2. In addition to class imbalance, we consider settings where we randomly remove training samples (left graph). We show this setting to determine that the accuracy gain/loss from output feature preservation is dependent on the available information, not the number of samples. Our experiments highlight both advantages and disadvantages of preserving features in the output representation: if the target distribution is sufficiently similar to the input distribution, additional preserved features correlate with the generalization objective and results in a performance increase. This can be seen from the accuracy improvement with spectral normalization under low imbalance severities or when samples are removed randomly. The opposite can be observed where, in contrast to random sample removal (left graph), we explicitly remove training and test samples to imbalance classes (right graph). Here, the training set contains more significant amounts of irrelevant information and spectral normalization significantly decreases the generalization performance.
@@ 4.2 Algorithm
Our method consists of three components: a principal feed-forward network, a constellation of shallow-deep network exits with individual internal classifiers, and a combination head (Figure 3). During training, the shallow-deep network exits are trained jointly with the feed-forward component, while the combination head is fitted after optimization on a validation set extracted from the training data $\mathcal{X}_{ID}$. We emphasize that our method does not require out-of-distribution validation samples as the combination head is fitted with data from the training set. During inference, the input sample traverses both the feed-forward network, as well as the shallow exits. The final output prediction is derived from the main network, while the uncertainty score is derived from a combination of the intermediate output logits. In addition to our description in the main paper, we provide implementation details and algorithm pseudo code in Appendix D.
@@
Shallow-Deep networks (SDNs) (Kaya et al., 2019) were originally introduced in the context of computation reduction. A SDN is a modified version of a conventional DNN where additional internal classifiers are placed on intermediate representations to produce preliminary predictions. In our context, we utilize internal classifiers to produce intermediate logits for uncertainty estimation. Formally, given the intermediate layer $l$ of the principal feed-forward network with the latent representation $h^{l}_{w_{l}}(\mathbf{x})\in\mathbb{R}^{M_{l}}$, an internal classifier is a shallow network $f^{l}_{w_{l}}:\mathbb{R}^{M_{l}}\rightarrow\mathbb{R}^{K}$ that produces prediction logits for $K$ classes. Regarding positioning, we place internal classifiers uniformly across the feed forward network (Kaya et al., 2019).
@@ 7 Discussion and Limitations
A central observation we made in this work is that enforcing feature preservation by constraining model representations can be harmful to the model performance. We highlight application relevance as a key requirement for effective representation constraints. In practice, the characteristic can be undesirable as training distributions can severely differ from deployment. We propose single-pass transitional feature preservation through intermediate representations to address these disadvantages. While our approach is effective, we do not claim that our improvements solve feature preservation in uncertainty estimation entirely. In particular, we propose one instance of TFP through SDNs that comes with its own set of limitations: similar to iterative methods, the success depends on the amount of source representations in $\Delta h$ to preserve features. For this purpose, SDNs are less effective on shallow architectures with fewer intermediate options to extract from. Further, we acknowledge that implementing internal classifiers comes with a large set of hyperparameters increasing space complexity. In our implementation, we use standard settings that show promising performance without significantly impacting the space complexity. However, other applications may require more elaborate hyperparameter explorations and additional internal classifiers. For this reason, we see reducing the space complexity of internal classifiers or the development of efficient hyperparameter tuners as promising research directions. Finally, we chose the combination of SDNs with Gaussian Processes out of simplicity and the strong empirical performance. However, key novelties of this paper (SDNs and singe-pass TFP) are not limited to one uncertainty score (Gaussian Processes) and are building blocks for single-pass uncertainty methods. We encourage researchers to implement different combinations of single-pass TFP and uncertainty scores to advance the field of single-pass uncertainty estimation.